OpenStudy (pjones14):

A tire has a pressure of 780 mmHg at 27 C. If the tire is heated to 47 C on hot pavement, the new pressure will be 310 mmHg 280 mmHg 830 mmHg 590 mmHg None of the above

OpenStudy (prathamesh_m):

use gay-lussac's law: pressure times temperature is constant for a given volume.

OpenStudy (australopithecus):

If you take the ideal gas equation: PV = nRT assuming constant amount of gas and volume you have two equations based on whats given: Eq. 1 \[\frac{P_1}{T_1} = \frac{nR}{V}\] Eq. 2 \[\frac{P_2}{T_2} = \frac{nR}{V}\] Since n, R and V are constant you get: \[\frac{P_1}{T_1} = \frac{P_2}{T_2}\] Or essentially the law that the person above stated

OpenStudy (jfraser):

just be sure to convert your temperatures into Kelvin before using the gas law @Australopithecus has given